# HGSLoc: 3DGS-based Heuristic Camera Pose Refinement

### v2

@@ Abstract
Visual localization refers to the process of determining camera poses and orientation within a known scene representation. This task is often complicated by factors such as illumination changes and variations in viewing angles. In this paper, we propose HGSLoc, a novel lightweight, plug-and-play pose optimization framework, which integrates 3D reconstruction with a heuristic refinement strategy to achieve higher pose estimation accuracy. Specifically, we introduce an explicit geometric map for 3D representation and high-fidelity rendering, allowing the generation of high-quality synthesized views to support accurate visual localization. Our method demonstrates a faster rendering speed and higher localization accuracy compared to NeRF-based neural rendering localization approaches. We introduce a heuristic refinement strategy, its efficient optimization capability can quickly locate the target node, while we set the step-level optimization step to enhance the pose accuracy in the scenarios with small errors. With carefully designed heuristic functions, it offers efficient optimization capabilities, enabling rapid error reduction in rough localization estimations. Our method mitigates the dependence on complex neural network models while demonstrating improved robustness against noise and higher localization accuracy in challenging environments, as compared to neural network joint optimization strategies. The optimization framework proposed in this paper introduces novel approaches to visual localization by integrating the advantages of 3D reconstruction and heuristic refinement strategy, which demonstrates strong performance across multiple benchmark datasets, including 7Scenes and DB dataset. The implementation of our method will be made open-source.
@@ I INTRODUCTION
Visual localization is a research direction aimed to determine the pose and orientation of a camera within a known scene by analyzing and processing image data. This technique has significant applications in various fields, such as augmented reality (AR), robot navigation, and autonomous driving. By enabling devices to accurately identify their spatial location in complex 3D environments, visual localization facilitates autonomous navigation, environmental awareness, and real-time interaction. The core objective of visual localization is to estimate the camera’s absolute pose. However, this task is challenging due to factors like illumination changes, dynamic occlusions, and variations in viewing angles, necessitating the development of robust and efficient algorithms to address these complexities.
Two major categories of methods in visual localization are Absolute Pose Regression (APR)[1, 2, 3, 4, 5, 6, 7, 8] and Scene Coordinate Regression (SCR)[9, 10, 11]. APR is an end-to-end deep learning approach that directly regresses the camera’s pos from the input image. The key advantages of APR lie in its simplicity and computational efficiency. However, APR exhibits notable limitations, particularly in complex or previously unseen environments, where its generalization capability is weak[12]. In contrast, SCR adopts an indirect strategy for pose estimation. It first predicts the 3D scene coordinates of each image pixel using a deep learning model, followed by the computation of the camera’s pose through spatial transformation of these coordinates. While SCR demonstrates high accuracy and robustness in familiar scenes, it incurs substantial computational costs due to the need to predict a large number of pixel-wise coordinates.
@@
In this paper, we propose a novel paradigm based on classical visual localization methods, aimed at improving the precision and accuracy of pos estimation in visual localization by integrating 3D reconstruction. Neural Radiance Field (NeRF)[13], a neural network-based 3D scene modeling approach, is capable of synthesizing and rendering high-quality 3D scene images through neural network training. However, NeRF’s pixel-wise training and inference mechanism results in significant computational overhead, limiting its practical applications. In contrast, 3D Gaussian Splatting (3DGS)[14] mitigates this issue by representing scene points as Gaussian distributions, thereby significantly reducing the data processing load during rendering. Furthermore, 3DGS leverages CUDA kernel functions to accelerate training and inference, making it a prominent method in the field of 3D reconstruction. In known or partially known static environments, several approaches, such as 3DGS-ReLoc[15] and GSLoc[16], have been developed. The 3DGS-ReLoc method requires grid search for efficiency in coarse localization using the normalized cross-correlation (NCC)[17] metric, which affects the localization accuracy.The GSLoc method has more steps and also uses MASt3R[18] for assisted localization. Whereas, our method is a lightweight framework that enables efficient positional optimization for any image. As shown in Fig. 1, by incorporating 3DGS, richer geometric information is available for pose estimation, and through heuristic optimization of coarse pos estimates, the accuracy of localization can be significantly enhanced in complex scenes.
Absolute Pose Regression (APR) and Scene Coordinate Regression (SCR) provide coarse pose estimates that serve as a foundation for further refinement. To achieve high-quality scene rendering, we introduce the 3D Gaussian Splatting (3DGS), which enriches the database imagery by constructing a dense point cloud, facilitating more detailed scene reconstruction. Building on this, we employ a heuristic refinement algorithm[19] to optimize the estimated poses. With its efficient pathfinding capabilities, combined with a custom-designed heuristic function, the algorithm efficiently adjusts the rendered view of the current pose to match the query image, resulting in more precise pose alignment. Our modular approach significantly reduces dependence on expensive neural network training, offering a more cost-effective solution compared to deep learning methods typically used for pose optimization. Additionally, our method exhibits strong generalization capabilities, maintaining rapid convergence and substantial improvements in pose accuracy, even in the presence of noisy pose data. This adaptability is particularly valuable in practical applications, as it ensures that the proposed method can be deployed across diverse platforms and data quality levels, providing a robust solution for a wide range of scenarios. The effectiveness of our approach is demonstrated through experiments conducted on several benchmark datasets, including 7Scenes and DB. These results underscore the method’s performance on classical visual localization datasets as well as those related to 3D Gaussian splatting. The contributions of our approach are summarized as follows:
@@
  • Our proposed framework achieves higher localization accuracy than NeRF-based neural rendering localization approaches [20] and outperforms neural network joint pose optimization strategy in noisy conditions.
@@ II RELATED WORK
In this section, we introduce visual localization methods and 3D Gaussian Splatiing.
@@ II-A Visual localization
PoseNet represents a foundational work in the domain of Absolute Pose Regression (APR)[1, 2, 3, 4, 5, 6, 7, 8], pioneering the direct regression of pose from image data using convolutional neural networks (CNNs). Unlike traditional localization techniques, which typically involve intricate feature extraction, matching, and geometric computation, PoseNet[1] introduces an end-to-end framework that seamlessly integrates these steps into a unified neural network learning process. This approach simplifies the mapping of image data to pose estimation, making it highly suitable for visual localization tasks across diverse environments. Building on PoseNet, MS-Transformer[7] enhances performance by incorporating global context modeling, enabling more effective handling of objects and structures at various scales within an image. The introduction of a multi-head self-attention mechanism allows for a better understanding of complex scenes, leading to significant improvements in pose regression accuracy. Likewise, DFNet[6] extends the capabilities of APR by integrating information from multimodal sensors, offering more comprehensive and detailed modeling of visual scenes. This fusion of multimodal data leverages the complementary strengths of different data sources, enhancing robustness and adaptability to various environmental factors. However, despite the advantages offered by APR methods, they remain vulnerable to noise and environmental variability. Under adverse conditions, such as poor lighting, unfavorable weather, or occlusions, the regression models’ accuracy in pose estimation can degrade significantly.
Scene Coordinate Regression (SCR) methods[9, 10, 11] estimate camera pose by learning the mapping between image pixels and corresponding 3D scene coordinates. These approaches bypass the complex feature matching procedures characteristic of traditional localization methods, thereby enhancing the efficiency and robustness of pose estimation. DSAC*[9] further advances SCR by introducing a differentiable hypothesis selection mechanism, allowing the model to learn how to choose the optimal pose hypothesis during network training. Additionally, it accommodates both RGB and RGB-D image inputs, incorporating depth map information into the pose estimation process, which enhances the model’s ability to interpret and manage complex scenes. On the other hand, ACE[10] accelerates feature matching by optimizing the encoding and decoding of image coordinates, which enables faster processing. Furthermore, it demonstrates resilience to noise and lighting variations, improving its robustness in dynamic or less controlled environments. By addressing these common challenges, ACE contributes to more reliable pose estimation in scenes where traditional methods may struggle to maintain accuracy.
@@
3D Gaussian Splatting (3DGS)[14], an emerging method in 3D reconstruction, has rapidly gained prominence since its introduction. This method significantly accelerates the synthesis of new views by modeling the scene with Gaussian ellipsoids and utilizing advanced rendering methods. Within the realm of 3DGS research, various techniques have enhanced and optimized 3DGS in different aspects, such as quality improvement[21], compression and regularization[22], dynamic 3D reconstruction[23], and handling challenging inputs[24]. The advancement of 3DGS methods not only enhances the quality of scene reconstruction but also speeds up rendering, offering novel and improved approaches for visual localization tasks. For instance, GSLoc leverages rendered images from new viewpoints for matching and pose optimization, while the InstantSplat[25] method, utilizing DUSt3R[26], achieves rapid and high-quality scene reconstruction by jointly optimizing poses with 3D Gaussian parameters. Our proposed method builds upon 3DGS reconstructed scenes and employs heuristic pose optimization to enhance pose accuracy in specific scenarios while preserving the original pose accuracy.
@@ III-B Heuristic Algorithm Implementation
Heuristic approaches[19] are often implemented to path planning and graph search that combines the strengths of depth-first search (DFS) and breadth-first search (BFS). It has been widely applied to various real-world problems, including game development, robot navigation, and geographic information systems (GIS). The primary goal of the heuristic algorithm is to efficiently find the optimal path from an initial node to a goal node, where each node represents a state within the search space. The algorithm relies on an evaluation function, $f(n)$, to prioritize nodes for expansion. This function typically consists of two components:
@@
Where the $I_{q}$ represents the current query image and ${I_{n}}_{child}$ represents the rendering image of current child node.
The heuristic function effectively guides the algorithm toward the optimal pose, ultimately identifying the pose that produces a rendered image most similar to the query image. We provide the pseudo-code for the algorithm’s implementation in Tab. I. In this pseudo-code, OpenList is used to store nodes awaiting expansion, while ClosedList contains nodes that have already been expanded.
@@ IV Experiment
NeRFMatch[28] 0.9/0.3 1.3/0.4 1.6/1.0 3.3/0.7 3.2/0.6 1.3/0.3 7.2/1.3 2.7/0.70
Marepo+HGSLoc 1.5/0.68 1.4/0.62 1.5/0.92 2.7/0.80 1.8/0.46 2.2/0.63 4.8/1.34 2.3/0.78
@@ IV-A Implementation
The deep learning framework employed in this work is PyTorch[29]. Each scene is reconstructed using 3D Gaussian Splatting (3DGS) with 30,000 training iterations, running on RTX 4090 GPUs. For the 7Scenes datasets, we adopt the SfM ground truth (GT) provided by [30].
@@ IV-B Datasets, Metrics and Baselines
We evaluated our method on two public datasets: 7scenes and Deep Blending. In the case of the 7scenes datasets[31, 32], the official test lists were used as query images, while the remaining images were utilized for training. For the Deep Blending dataset, we specifically selected the drjohnson and playroom scenes, and we constructed a test image set following the 1-out-of-8 approach suggested by Mip-NeRF[33].
@@
We show the median rotation and translation error, and also provide the ratio of pose error within 1cm/1°.
@@
Our approach builds on an initial coarse pose estimation. For the APR[1, 2, 3, 4, 5, 6, 7, 8] framework, we have selected the widely recognized Marepo[8] method as the benchmark for comparison. Similarly, for the SCR[9, 10, 11] framework, we have chosen the classical ACE[10] method as the benchmark for comparison.
@@ IV-C Analysis of results
       NRP        Marepo+HGSLoc        19.1
       NRP        ACE+HGSLoc        59.1
DB dataset
We selected two scenes, ”playroom” and ”drjohnson,” for testing. For both the Marepo[8] and ACE[10] methods, we observed that the coarse pose errors were significantly large. This may be attributed to the higher complexity of the DB dataset compared to the 7Scenes datasets, as well as the limited training data, which may have prevented model convergence. Consequently, we utilized an alternative method (HLoc) that leverages point clouds to obtain an initial pose estimate and compared the results. As shown in Tab. IV, the improvement from boosting is not pronounced, likely due to the high image quality of the DB dataset, which already provided relatively accurate preliminary poses with the HLoc framework. To better demonstrate the effectiveness of our pose optimization method, Tab. V introduces various levels of step noise, making the visualization results more intuitive.
@@
TABLE V: We show the median translation and rotation error (m/°) for the poses with noise and for the poses after optimization. (q2, t1) denotes the introduction of noise at the percentile of qvec, decile of tvec, and the rest is the same.
@@
As shown in Tab. VI, to further demonstrate the effectiveness of our method, we compare it with an alternative joint optimization strategy[25]. For this comparison, a noise level of $1\times 10^{-3}$ granularity is introduced to the initial pose. Our method employs heuristic optimization based on high-quality scene reconstruction obtained through the 3DGS[14] method, whereas the alternative strategy jointly optimizes both the scene reconstruction and the initial pose[25].
TABLE VI: We show the median translation and rotation error (m/°) for heuristic optimization and joint optimization strategies.
@@
By inputting the pose into the 3D reconstructed scene, we generate a rendered image that visualizes the pose. Each query image corresponds to the GT pose, and the discrepancy between the estimated pose and the GT pose is reflected in the rendered images from various viewpoints. To better observe this error and the improvement achieved through our optimization method, we select viewpoints with significant accuracy improvements for qualitative analysis. Fig. 3 demonstrate that, when using our framework on the 7Scenes datasets, the rendered images more closely match the GT images. Fig. 4illustrates the results of applying our framework to noisy poses in the DB dataset, showing that our method effectively refines the original pose, resulting in rendered images that closely resemble the GT images.
@@
Figure 4: Each subimage is divided by a diagonal line, with the image rendered by the estimated pose on the lower left and the GT image on the upper right. The diagonal lines in the optimized comparison image appear less distinct, reflecting improved alignment with the GT image. HGSLoc demonstrates its effectiveness in refining pose estimation, achieving precise values while mitigating the impact of band noise.
@@
In our method, we use the sum of pixel-by-pixel differences as the heuristic function. To demonstrate the effectiveness of this heuristic function, Tab. VII compares the results obtained using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) as alternative heuristic functions. Higher values of PSNR and SSIM indicate better image quality and structural similarity, whereas we would like to see them take the opposite number as the value of the heuristic function is as small as possible. To illustrate the impact of different heuristic functions more clearly, we applied these comparisons to the DB dataset, which introduces significant noise.
@@
TABLE VII: We show the median translation and rotation error (m/°) for poses with noise and for poses after optimization using different heuristic functions.
@@
V CONCLUSIONS
In this study, we propose a lightweight, plug-and-play visual localization optimization framework that combines heuristic refinement strategy with 3D reconstruction to significantly enhance pose estimation accuracy, achieving SOTA performance on two datasets. Compared to NeRF-based neural rendering localization methods[20], the proposed approach demonstrates superior rendering speed and enhanced localization accuracy. Through the integration of well-designed heuristic functions, the method efficiently optimizes and rapidly reduces errors in coarse localization estimations. Our modular approach not only reduces reliance on complex neural network training, enhancing the algorithm’s flexibility and practicality, but also demonstrates robust performance in noisy environments, facilitating rapid convergence and higher accuracy. This robustness ensures that the method performs consistently across various platforms and data qualities. In summary, the integration of heuristic refinement strategy with 3D Gaussian distribution offers a novel and effective solution for visual localization, providing a valuable reference for the development and optimization of future visual localization systems.

### v3

@@ Abstract
Visual localization refers to the process of determining camera poses and orientation within a known scene representation. This task is often complicated by factors such as changes in illumination and variations in viewing angles. In this paper, we propose HGSLoc, a novel lightweight plug-and-play pose optimization framework, which integrates 3D reconstruction with a heuristic refinement strategy to achieve higher pose estimation accuracy. Specifically, we introduce an explicit geometric map for 3D representation and high-fidelity rendering, allowing the generation of high-quality synthesized views to support accurate visual localization. Our method demonstrates higher localization accuracy compared to NeRF-based neural rendering localization approaches. We introduce a heuristic refinement strategy, its efficient optimization capability can quickly locate the target node, while we set the step-level optimization step to enhance the pose accuracy in the scenarios with small errors. With carefully designed heuristic functions, it offers efficient optimization capabilities, enabling rapid error reduction in rough localization estimations. Our method mitigates the dependence on complex neural network models while demonstrating improved robustness against noise and higher localization accuracy in challenging environments, as compared to neural network joint optimization strategies. The optimization framework proposed in this paper introduces novel approaches to visual localization by integrating the advantages of 3D reconstruction and the heuristic refinement strategy, which demonstrates strong performance across multiple benchmark datasets, including 7Scenes and Deep Blending dataset. The implementation of our method has been released at https://github.com/anchang699/HGSLoc.
@@ I INTRODUCTION
Visual localization is a research direction that aims to determine the pose and orientation of a camera within a known scene by analyzing and processing image data. This technique has significant applications in various fields, such as augmented reality (AR), robotic navigation, and autonomous driving. By enabling devices to accurately identify their spatial location in complex 3D environments, visual localization facilitates autonomous navigation, environmental awareness, and real-time interaction. The core objective of visual localization is to estimate the camera’s absolute pose. However, this task is challenging due to factors such as illumination changes, dynamic occlusions, and variations in viewing angles, necessitating the development of robust and efficient algorithms to address these complexities.
Two major categories of visual location methods are Absolute Pose Regression (APR)[1, 2, 3, 4, 5, 6, 7, 8] and Scene Coordinate Regression (SCR)[9, 10, 11]. APR is an end-to-end deep learning approach that directly regresses the camera’s pose from the input image. The key advantages of APR lie in its simplicity and computational efficiency. However, APR exhibits notable limitations, particularly in complex or previously unseen environments, where its generalization capability is weak[12]. In contrast, SCR adopts an indirect strategy for pose estimation. It first predicts the 3D scene coordinates of each image pixel using a deep learning model, followed by the computation of the camera pose through the spatial transformation of these coordinates. Although SCR demonstrates high accuracy and robustness in familiar scenes, it incurs substantial computational costs due to the need to predict a large number of pixel-wise coordinates.
@@
In this paper, we propose a novel visual localization paradigm that enhances pose estimation accuracy through the integration of 3D reconstruction. Neural Radiance Field (NeRF)[13] is capable of synthesizing and rendering high-quality 3D scene images through neural network training. Some existing NeRF-based visual localization methods[14, 15] have gained widespread recognition. However, NeRF’s pixel-wise training and inference mechanism results in significant computational costs, limiting its practical applications. In contrast, 3D Gaussian Splatting (3DGS)[16] represents scene points as Gaussian distributions, significantly reducing computational overhead and leveraging CUDA acceleration for efficient training and inference. In known or partially known static environments, several approaches, such as 3DGS-ReLoc[17] and GSLoc[18], have been developed. The 3DGS-ReLoc method relies on grid search with the normalized cross-correlation (NCC)[19], which affects the localization accuracy. The GSLoc method requires several rounds of iterations to achieve the desired result in the case of poor initial pose estimation, and each iteration utilizes MASt3R[20] for assisted localization. Whereas, our method is a lightweight framework that enables efficient positional optimization for any image. As shown in Fig. 1, by incorporating 3DGS, richer geometric information is available for pose estimation, and through heuristic optimization of coarse pose estimates, the accuracy of localization can be significantly enhanced in complex scenes.
Absolute Pose Regression (APR) and Scene Coordinate Regression (SCR) provide coarse pose estimates for further refinement. To enhance scene rendering, we introduce 3D Gaussian Splatting (3DGS), which constructs a dense point cloud for high-fidelity reconstruction. Building on this, we employ a heuristic refinement algorithm[21] that efficiently adjusts the rendered view to better align with the query image, improving pose accuracy. This modular approach reduces reliance on computationally expensive neural network training, offering a more efficient alternative to deep learning-based pose optimization. Our method demonstrates strong generalization capabilities, maintaining high accuracy even in the presence of noisy pose data, making it adaptable across various platforms. Experimental results on benchmark datasets, including 7Scenes and Deep Blending, validate its effectiveness in visual localization tasks. The contributions of our approach are summarized as follows:
@@
  • Our proposed framework achieves higher localization accuracy than NeRF-based neural rendering localization approach [22] and outperforms neural network joint pose optimization strategy in noisy conditions.
@@ II RELATED WORK
In this section, we introduce visual localization methods and 3D Gaussian Splatting.
@@ II-A Visual localization
PoseNet pioneered Absolute Pose Regression (APR)[1, 2, 3, 4, 5, 6, 7, 8] by employing a convolutional neural network (CNN) to directly regress camera pose from images, bypassing traditional feature extraction and geometric computations. This end-to-end approach simplifies visual localization across diverse environments. MS-Transformer[7] improves APR by incorporating global context modeling and multi-head self-attention, enhancing scene understanding and pose accuracy. DFNet[6] further extends APR by integrating multimodal sensor data, increasing robustness. However, APR methods remain sensitive to noise and environmental variability, with accuracy degrading under poor lighting, adverse weather, or occlusions.
Scene Coordinate Regression (SCR)[9, 10, 11] estimates camera pose by mapping image pixels to 3D scene coordinates, eliminating the need for complex feature matching and improving efficiency. DSAC*[9] enhances SCR with a differentiable hypothesis selection mechanism and supports both RGB and RGB-D inputs, leveraging depth information for improved scene interpretation. ACE[10] optimizes image coordinate encoding and decoding to accelerate feature matching while enhancing robustness to noise and lighting variations, ensuring reliable pose estimation in challenging environments.
@@
3D Gaussian Splatting (3DGS)[16], an emerging method in 3D reconstruction, has rapidly gained prominence since its introduction. This method significantly accelerates the synthesis of new views by modeling the scene with Gaussian ellipsoids and utilizing advanced rendering methods. Within the realm of 3DGS research, various techniques have enhanced and optimized 3DGS in different aspects, such as quality improvement[23], compression and regularization[24], navigation[25], dynamic 3D reconstruction[26], and handling challenging inputs[27]. The advancement of 3DGS methods not only enhances the quality of scene reconstruction but also speeds up rendering, offering novel and improved approaches for visual localization tasks. For instance, GSLoc[18] leverages rendered images from new viewpoints for matching and pose optimization, while the InstantSplat[28] method, utilizing DUSt3R[29], achieves rapid and high-quality scene reconstruction by jointly optimizing poses with 3D Gaussian parameters. Our proposed method builds upon 3DGS reconstructed scenes and employs heuristic pose optimization to enhance pose accuracy in specific scenarios while preserving the original pose accuracy.
@@ III-B Heuristic Algorithm Implementation
Heuristic approaches[21] are often implemented to path planning and graph search, combining the strengths of depth-first search (DFS) and breadth-first search (BFS). It has been widely applied to various real-world problems, including game development, robotic navigation, and geographic information systems (GIS). The primary goal of the heuristic algorithm is to efficiently find the optimal path from an initial node to a goal node, where each node represents a state within the search space. The algorithm relies on an evaluation function, $f(n)$, to prioritize nodes for expansion. This function typically consists of two components:
@@
Where $I_{q}$ represents the current query image and ${I_{n}}_{child}$ represents the rendering image of current child node.
The heuristic function effectively guides the algorithm toward the optimal pose, ultimately identifying the pose that produces a rendered image most similar to the query image. We provide the pseudo-code for the algorithm’s implementation in Tab. I. In this pseudo-code, OpenList is used to store nodes awaiting expansion, while CloseList contains nodes that have already been expanded.
@@ IV Experiment
NeRFMatch[22] 0.9/0.3 1.1/0.4 1.5/1.0 3.0/0.8 2.2/0.6 1.0/0.3 10.1/0.7 2.8/0.73
Marepo+HGSLoc 0.7/0.33 1.4/0.62 1.5/0.92 2.2/0.70 1.8/0.46 2.2/0.63 4.8/1.34 2.1/0.71
@@ IV-A Implementation
The deep learning framework employed in this work is PyTorch[31]. Each scene is reconstructed using 3D Gaussian Splatting (3DGS) with 30,000 training iterations, running on RTX 4090 GPUs. For the 7Scenes dataset, we adopt the SfM ground truth (GT) provided by [32].
@@ IV-B Datasets, Metrics and Baselines
We evaluated our method on two public datasets: 7Scenes and Deep Blending. In the case of the 7Scenes datasets[33, 34], the official test lists were used as query images, while the remaining images were utilized for training. For the Deep Blending dataset[35], we selected four scenes and constructed a test image set following the 1-out-of-8 approach suggested by Mip-NeRF[36].
@@
We show the median rotation and translation errors, and also provide the ratio of pose error within 1cm/1°.
@@
Our approach is built upon an initial coarse pose estimation. For the APR[1, 2, 3, 4, 5, 6, 7, 8] framework, we have selected the widely recognized Marepo[8] method as the benchmark for comparison. Similarly, for the SCR[9, 10, 11] framework, we have chosen the classical ACE[10] method as the benchmark for comparison.
@@ IV-C Analysis of results
       NRP        Marepo+HGSLoc        25.8
       NRP        ACE+HGSLoc        59.2
Deep Blending dataset
We selected two indoor scenes, ”Playroom” and ”DrJohnson”, and two outdoor scenes, ”Boats” and ”NightSnow”, for testing. For both the Marepo[8] and ACE[10] methods, we observed that the coarse pose errors were significantly large. This may be attributed to the higher complexity of the Deep Blending dataset compared to the 7Scenes datasets, as well as the limited training data, which may have prevented model convergence. Consequently, we utilized an alternative method (HLoc[37]) that leverages point clouds to obtain an initial pose estimate and compared the results. As shown in Tab. IV, the improvement from boosting is not pronounced, likely due to the high image quality of the Deep Blending dataset, which already provided relatively accurate preliminary poses with the HLoc framework. To better demonstrate the effectiveness of our pose optimization method, Tab. V introduces various levels of step noise, making the visualization results more intuitive.
@@
        Boats         0.5/0.016         0.4/0.013
        NightSnow         0.2/0.024         0.2/0.019
TABLE V: We show the median translation and rotation errors (m/°) for the poses with noise and for the poses after optimization. (q2, t1) denotes the introduction of noise at the percentile of qvec, decile of tvec, and the rest is the same.
@@
(c) Boats
noise error refine error tvec↑ qvec↑
q2, t1 0.62/6.41 0.01/0.02 98.4% 99.7%
q2, t2 0.26/9.17 0.17/2.46 34.6% 73.2%
q3, t3 0.04/0.71 0.01/0.15 75.0% 78.9%
(d) NightSnow
noise error refine error tvec↑ qvec↑
q2, t1 0.49/7.79 0.01/0.03 98.0% 99.6%
q2, t2 0.29/7.62 0.07/0.58 75.9% 92.4%
q3, t3 0.03/0.69 0.01/0.10 66.7% 85.5%
As shown in Tab. VI, to further demonstrate the effectiveness of our method, we compare it with an alternative joint optimization strategy[28]. For this comparison, a noise level of $1\times 10^{-3}$ granularity is introduced to the initial pose. Our method employs heuristic optimization based on high-quality scene reconstruction obtained through the 3DGS[16] method, whereas the alternative strategy jointly optimizes both the scene reconstruction and the initial noisy pose[28].
TABLE VI: We show the median translation and rotation errors (m/°) for heuristic optimization and joint optimization strategies.
@@
By inputting the pose into the 3D reconstructed scene, we generate a rendered image that visualizes the pose. Each query image corresponds to the GT pose, and the discrepancy between the estimated pose and the GT pose is reflected in the rendered images from various viewpoints. To better analyze errors and optimization improvements, we select viewpoints where significant accuracy gains are observed. Fig. 3 demonstrate that, when using our framework on the 7Scenes datasets, the rendered images more closely match the GT images. Fig. 4 illustrates the results of applying our framework to noisy poses in the Deep Blending dataset, showing that our method effectively refines the original pose, resulting in rendered images that closely resemble the GT images. However, experiments on the Cambridge Landmarks dataset[1] yielded suboptimal results, likely due to its limited viewpoint coverage, which leads to gaps in the 3D reconstruction of the scenes. In contrast, the 360° closed-loop image structure of the Deep Blending dataset enhance the effectiveness of pose optimization.
@@
(a) Indoor Scenes Refer to caption (b) Outdoor Scenes
Figure 4: Each subimage is divided by a diagonal line, with the image rendered by the estimated pose on the lower left and the GT image on the upper right. The diagonal lines in the optimized comparison image appear less distinct, reflecting improved alignment with the GT image. HGSLoc demonstrates its effectiveness in refining pose estimation, achieving precise values while mitigating the impact of noise.
@@
In our method, we use the sum of pixel-by-pixel differences as the heuristic function. To demonstrate the effectiveness of this heuristic function, Tab. VII compares the results obtained using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) as alternative heuristic functions. Higher values of PSNR and SSIM indicate better image quality and structural similarity, whereas we would like to see them take the opposite number as the value of the heuristic function is as small as possible. To illustrate the impact of different heuristic functions more clearly, we applied these comparisons to the Deep Blending dataset, which introduces a relatively large level of noise.
@@
TABLE VII: We show the median translation and rotation errors (m/°) for poses with noise and for poses after optimization using different heuristic functions.
@@
Boats 0.62/6.41 0.01/0.02 0.44/3.93 0.59/5.97
NightSnow 0.49/7.79 0.01/0.03 0.49/7.03 0.46/7.25
V LIMITATIONS AND FUTURE WORK
While our method successfully achieves the desired results, it still has certain limitations. During the optimization process, typically 300–400 nodes need to be expanded, with multiple Gaussian renderings performed for each expansion, which significantly constrains the method’s real-time performance. Future work will focus on optimizing pose representation to reduce the number of Gaussian renderings required per node expansion, thereby improving computational efficiency and enhancing the method’s applicability to real-time scenarios.
VI CONCLUSIONS
This study introduces a lightweight, plug-and-play visual localization framework that integrates heuristic refinement with 3D reconstruction to enhance pose accuracy, achieving state-of-the-art performance on two datasets. Compared to NeRF-based methods[22], it achieves superior localization accuracy by efficiently refining coarse estimations using heuristic functions. The modular approach not only reduces dependence on complex neural network training, but also demonstrates robust performance in noisy environments. By combining heuristic refinement with a 3D Gaussian distribution, this approach offers a novel and effective solution, providing valuable reference for the future development of visual localization systems.